# Agentic Design of Compositional Machines

### v1

@@ Abstract
The design of complex machines stands as both a marker of human intelligence and a foundation of engineering practice. Given recent advances in large language models (LLMs), we ask whether they, too, can learn to create. We approach this question through the lens of compositional machine design: a task in which machines are assembled from standardized components to meet functional demands like locomotion or manipulation in a simulated physical environment. To support this investigation, we introduce BesiegeField, a testbed built on the machine-building game Besiege, which enables part-based construction, physical simulation and reward-driven evaluation. Using BesiegeField, we benchmark state-of-the-art LLMs with agentic workflows and identify key capabilities required for success, including spatial reasoning, strategic assembly, and instruction-following. As current open-source models fall short, we explore reinforcement learning (RL) as a path to improvement: we curate a cold-start dataset, conduct RL finetuning experiments, and highlight open challenges at the intersection of language, machine design, and physical reasoning.
@@ 3 BesiegeField: Playground for Compositional Machine Design
To this end, we create BesiegeField, an interactive environment adapted from the machine-building game Besiege, in which players design medieval machines to complete tasks such as destroying castles. Powered by the built-in physics engine, BesiegeField supports physical simulation of mechanical systems such as vehicles and catapults in user-customized environments with terrains, obstacles, external forces (e.g., wind and gravity), and co-existing agents. The environment provides nearly 80 types of building blocks (examples illustrated in Fig. 10), including passive ones like drills and logs, and powered ones like powered cogs and wheels. Machines are constructed by sequentially attaching new parts to vacant and attachable faces of existing blocks, starting from a root block and thus forming a “construction tree” (indeed a directly acyclic graph (DAG), in the sense of operation orders; one block can has two parents in the DAG; the actual structures may contain loops). Powered blocks can receive control commands, allowing machines to be operated precisely. During simulation, complete state information (e.g., the position and velocity of each block in the constructed machine) can be recorded for model feedback. Finally, the environment supports custom modifications and can be extended with additional block types and richer physics (e.g., simple fluid simulation). Further details are explained in Appendix B.
BesiegeField is unique in balancing real-world geometry and physics, part-level semantics, and simple compositional rules. Block-stacking environments like LEGO (Fan et al., 2022) and Minecraft (Fan et al., 2022; Pun et al., 2025) allow intuitive combinatorial assembly but do not natively provide realistic physical simulation and rely on generic blocks with limited semantic meaning. CAD modeling (Li et al., 2025) captures fine-grained geometry and interactions, but its complexity makes rules cumbersome and sequences prohibitively long. By contrast, BesiegeField uses semantically meaningful parts with cuboid-like construction rules—supporting realistic physics while remaining abstract enough for tractable composition. This calibrated balance enables the study of compositional creativity and geometric reasoning at a level of difficulty that both differentiates algorithms and permits rapid experimentation. Moreover, unlike prior environments, BesiegeField supports machine destruction, adding durability and failure analysis to the design space.
@@ 4.1 Benchmark Settings
Machine representations. In BesiegeField, the default XML representation records all blocks with global 3D positions and uses a built-in algorithm to recover connections. Such a representation, however, does not well encode machine structures. Instead, we propose a parsimonious representation aligned with the game’s building logic, based on pairwise relative attachment relationships (i.e., how one part is rotated and attached to another). Details are explained in Appendix D.3.
Figure 3: Demonstration of the default XML representation and our construction tree representation. Parent block info is in blue and child info is in red.
@@ 4.3 Key Empirical Observations
Machine representation. We experiment with a coordinate-only representation derived from the default XML (Appendix D.1) and our construction tree representation. Results show that the coordinate-only representation performs significantly worse (Table 7), implying that explicit structural information is necessary for LLM understanding.
@@ 5.1 Experimental Settings
Cold-start finetuning and dataset curation. Following recent RLVR practices (Lambert et al., 2025; Yue et al., 2025; Zhu et al., 2025a), we curated a small dataset to cold-start LLMs by aligning their reasoning process with expert CoT. Specifically, we collected textual descriptions of machine functionalities from Besiege player communities and prompted Gemini 2.5 Pro to generate corresponding machines. After filtering out invalid generations, we obtained 9,984 valid machine–CoT pairs. We then used this dataset to perform supervised finetuning on Qwen-2.5-14B-Instruct for 12 epochs. Additional training details are provided in Appendix F.2.
@@ Acknowledgment
We sincerely thank the developers of Besiege for creating the game and fostering an open and vibrant community, without which our exploration of this exciting idea would not have been possible.
@@ Appendix B Details on the BesiegeField Environment
Figure 10: Common useful blocks in constructing standard machines in BesiegeField.
@@ B.3 Blocks
Out of the 75 construction blocks provided by Besiege, we filter out a list of 27 blocks (shown in Fig. 10) that are most relevant to build machines with classical mechanical mechanisms such as levers and trusses.
@@ Appendix C Search Strategies in Machine Modification Loops
Apart from the MCTS strategy used in the main experiments, we also evaluate two alternatives: (i) best-of-N, where we select the best-performing machine out of N candidates, and (ii) random search, which mimics best-of-N but instead selects a random candidate. For clarity, we refer to one consecutive “querier–refiner” call as a search node (consistent with our MCTS setup). Unlike classical MCTS or best-of-N, here each search node is allowed up to five retries to prevent child statistics from being too sparse. We perform $R$ search rounds, each aiming to obtain $5$ valid candidate machines (though this may fail; if fewer than 5 are found, the parent node’s machine is used as a candidate). Full algorithmic details are provided in Algorithm 1, Algorithm 2, and Algorithm 3. In Fig. 9 we show the improvement of machine performance with respect to the number of search rounds used. In Fig. 17 we compare the efficiency of different search methods in our agentic compositional machine design setting.
@@ D.1 Machine Representation
D.2 Global Position-based Representation
@@
D.3 Construction Tree Representation
@@ Appendix G Additional Ablation Studies
Additional RL results. In Fig. 23 and 25 and , we show the maximum scores achieved in the environments with different RL methods plus the validity rate of machines. We visualize the maximum score since, in the case when one is allowed to use inference-time scaling techniques, the best performing machines are the ones people care most about. We show that our settings with Pass@64 training achieves the best maximum score with two different random seeds. In additiona, in Fig. 27, we visualize the corresponding Best@N metrics.
For completeness, we also visualize the results with our default setting on the task car in Fig. 24, 26 and 28.

### v2

@@ Abstract
The design of complex machines stands as both a marker of human intelligence and a foundation of engineering practice. Given recent advances in large language models (LLMs), we ask whether they, too, can learn to create. We approach this question through the lens of compositional machine design: a task in which machines are assembled from standardized components to meet functional demands like locomotion or manipulation in a simulated physical environment. With this simplification, machine design is expressed as writing XML-like code that explicitly specifies pairwise part connections. To support this investigation, we introduce BesiegeField, a testbed built on the machine-building game Besiege, which enables part-based construction, physical simulation and reward-driven evaluation. Using BesiegeField, we benchmark state-of-the-art LLMs with agentic workflows and identify key capabilities required for success, including spatial reasoning, strategic assembly, and instruction-following. As current open-source models fall short, we explore reinforcement learning (RL) as a path to improvement: we curate a cold-start dataset, conduct RL finetuning experiments, and highlight open challenges at the intersection of language, machine design, and physical reasoning.
@@ 3 BesiegeField: Playground for Compositional Machine Design
To this end, we create BesiegeField, an interactive environment adapted from the machine-building game Besiege, in which players design medieval machines to complete tasks such as destroying castles. Powered by the built-in physics engine, BesiegeField supports physical simulation of mechanical systems such as vehicles and catapults in user-customized environments with terrains, obstacles, external forces (e.g., wind and gravity), and co-existing agents. The environment provides nearly 80 types of building blocks , including passive ones like drills and logs, and powered ones like powered cogs and wheels. Machines are constructed by sequentially attaching new parts to vacant and attachable faces of existing blocks, starting from a root block and thus forming a “construction tree” (indeed a directly acyclic graph (DAG), in the sense of operation orders; one block can has two parents in the DAG; the actual structures may contain loops). Powered blocks can receive control commands, allowing machines to be operated precisely. During simulation, complete state information (e.g., the position and velocity of each block in the constructed machine) can be recorded for model feedback. Finally, the environment supports custom modifications and can be extended with additional block types and richer physics (e.g., simple fluid simulation). Further details are explained in Appendix B.
BesiegeField is unique in balancing real-world geometry and physics, part-level semantics, and simple compositional rules. Block-stacking environments like LEGO (Fan et al., 2022) and Minecraft (Fan et al., 2022; Pun et al., 2025) allow intuitive combinatorial assembly but do not natively provide realistic physical simulation and rely on generic blocks with limited semantic meaning. CAD modeling (Li et al., 2025) captures fine-grained geometry and interactions, but its complexity makes rules cumbersome and sequences prohibitively long. By contrast, BesiegeField uses semantically meaningful parts with cuboid-like construction rules-supporting realistic physics while remaining abstract enough for tractable composition. This calibrated balance enables the study of compositional creativity and geometric reasoning at a level of difficulty that both differentiates algorithms and permits rapid experimentation. Moreover, unlike prior environments, BesiegeField supports machine destruction, adding durability and failure analysis to the design space.
@@ 4.1 Benchmark Settings
Machine representations. In BesiegeField, the default position-based representation records all blocks with global 3D positions and uses a built-in algorithm to recover connections. Such a representation, however, does not well encode machine structures. Instead, we propose a parsimonious representation aligned with the game’s building logic, based on pairwise relative attachment relationships (i.e., how one part is rotated and attached to another). Details are explained in Appendix D.1.2.
Figure 3: Demonstration of the default position-based representation and our construction tree representation. Parent block info is in blue and child info is in red.
@@ 4.3 Key Empirical Observations
Machine representation. We experiment with a coordinate-only representation derived from the default position-based (Appendix D.1) and our construction tree representation. Results show that the coordinate-only representation performs significantly worse (Table 7), implying that explicit structural information is necessary for LLM understanding.
@@ 5.1 Experimental Settings
Cold-start finetuning and dataset curation. Following recent RLVR practices (Lambert et al., 2025; Yue et al., 2025; Zhu et al., 2025a), we curated a small dataset to cold-start LLMs by aligning their reasoning process with expert CoT. Specifically, we collected textual descriptions of machine functionalities from Besiege player communities and prompted Gemini 2.5 Pro to generate corresponding machines. After filtering out invalid generations, we obtained 9,984 valid machine-CoT pairs. We then used this dataset to perform supervised finetuning on Qwen-2.5-14B-Instruct for 12 epochs. Additional training details are provided in Appendix F.2.
@@ Acknowledgment
We sincerely thank the developers of Besiege for creating such an inspiring game and for fostering an open, vibrant player community, without which our exploration of this idea would not have been possible. We also extend our gratitude to the developers of BepInEx, whose work made it possible for us to unlock the full potential of Besiege.
@@ B.3 Blocks
Out of the 75 construction blocks provided by Besiege, we filter out a list of 27 blocks that are most relevant to build machines with classical mechanical mechanisms such as levers and trusses.
@@ Appendix C Search Strategies in Machine Modification Loops
Apart from the MCTS strategy used in the main experiments, we also evaluate two alternatives: (i) best-of-N, where we select the best-performing machine out of N candidates, and (ii) random search, which mimics best-of-N but instead selects a random candidate. For clarity, we refer to one consecutive “querier-refiner” call as a search node (consistent with our MCTS setup). Unlike classical MCTS or best-of-N, here each search node is allowed up to five retries to prevent child statistics from being too sparse. We perform $R$ search rounds, each aiming to obtain $5$ valid candidate machines (though this may fail; if fewer than 5 are found, the parent node’s machine is used as a candidate). Full algorithmic details are provided in Algorithm 1, Algorithm 2, and Algorithm 3. In Fig. 9 we show the improvement of machine performance with respect to the number of search rounds used. In Fig. 16 we compare the efficiency of different search methods in our agentic compositional machine design setting.
@@ D.1 Machine Representation
D.1.1 Global Position-based Representation
@@
D.1.2 Construction Tree Representation
@@ Appendix G Additional Ablation Studies
Additional RL results. In Fig. 22 and 24 and , we show the maximum scores achieved in the environments with different RL methods plus the validity rate of machines. We visualize the maximum score since, in the case when one is allowed to use inference-time scaling techniques, the best performing machines are the ones people care most about. We show that our settings with Pass@64 training achieves the best maximum score with two different random seeds. In additiona, in Fig. 26, we visualize the corresponding Best@N metrics.
For completeness, we also visualize the results with our default setting on the task car in Fig. 23, 25 and 27.